OpenStudy (anonymous):

What is the simplified form of the expression? 7m2 + 6.5n – 4n + 2.5m2 – n

OpenStudy (anonymous):

did you mean \[7m^2+6.5n-4n+2.5m^2-n\]?

OpenStudy (anonymous):

combine common factors

OpenStudy (anonymous):

um yes thats what i meant

OpenStudy (anonymous):

9.5m^2+1.5m :)

OpenStudy (anonymous):

sorry 1.5n

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

im not that gud at algebra i need a tuture

OpenStudy (anonymous):

your gona learn here

OpenStudy (anonymous):

i hope so

OpenStudy (anonymous):

what is the order of \[\sqrt{5}, -0.1, \left(\begin{matrix}-5 \\ 3\end{matrix}\right), 0.7, \sqrt{2} from least \to greatest?\]